OpenStudy (anonymous):

someone please help i have no idea how to do this!! Choose the system of equations which matches the following graph. a line includes points 0 comma 4 and 2 comma 3. A line includes points 0 comma 3 and 4 comma 1 x – 2y = 8 2x + 4y = 12 x – 2y = 8 2x – 4y = 12 x + 2y = 8 2x + 4y = 12 x + 2y = 8 2x – 4y = 12

Directrix (directrix):

Is this --> "A line includes points 0 comma 4 and 2 comma 3. A line includes points 0 comma 3 and 4 comma 1" a description of the graph that is not shown?

OpenStudy (anonymous):

the graph is shown, yes

OpenStudy (anonymous):

that s the description, yeah

OpenStudy (anonymous):

please help me, im desperate! lol

Directrix (directrix):

Are the lines shown in the diagram parallel?

OpenStudy (anonymous):

yes

Directrix (directrix):

I'm talking about these two lines which only you can see: a line includes points 0 comma 4 and 2 comma 3. A line includes points 0 comma 3 and 4 comma 1

OpenStudy (anonymous):

yes, theyre parallel

Directrix (directrix):

Of the options, the correct one would have to be parallel lines. Parallel lines have the same slope but different y-intercepts in this case.

OpenStudy (anonymous):

so which of the options would it be?

Directrix (directrix):

We have to find the slopes.

OpenStudy (anonymous):

how do i do that

Directrix (directrix):

For each pair, change to slope-intercept form of line. For example the first pair of equations: 1)x – 2y = 8 2) x + 4y = 12

Directrix (directrix):

1) x – 2y = 8 x – 2y = 8 -2y = 8 - x y = x - 4 --> slope 1, y-intercept -4

Directrix (directrix):

2) x + 4y = 12 4y = 12 - x y = -(1/4) x + 3; slope -1/4 and y-intercept 3

Directrix (directrix):

The first pair of options is not the answer.

OpenStudy (anonymous):

would it be the last answer?

Directrix (directrix):

I am thinking that the third of fourth option is the one. Why do you think it is the last option?

Directrix (directrix):

On the last option, the slopes are not the same. Check.

Directrix (directrix):

The third option equations both have slope -1/2 and one has y-intercept 4 and the other 3. The third option is my choice for the correct answer.

Directrix (directrix):

What do you think?

OpenStudy (anonymous):

okay, i understand why its not the last one, thank u so much for your help!

Directrix (directrix):

Glad to help.